An air force plane flies at a speed that is 30 kilometers per hour faster than a cargo plane. The air force plane flies 500 km in the amount of time that the cargo plane flies 440 km. Find the speed of each plane.
Found 2 solutions by VFBundy, MathTherapy:
Answer by VFBundy(438)

Cargo plane
Distance = 400
Rate = r
Time = 400/r

Air force plane
Distance = 500
Rate = r + 30
Time = 500/(r + 30)

400/r = 500/(r + 30)

400(r + 30) = 500r

400r + 12000 = 500r

100r = 12000

r = 120

Cargo plane
Rate = r = 120

Air force plane
Rate = r + 30 = 150

Answer by MathTherapy(10855)


An air force plane flies at a speed that is 30 kilometers per hour faster than a cargo plane. The air force plane flies 500 km in the amount of time that the cargo plane flies 440 km. Find the speed of each plane.
Correct answers: